# Dyson Equation for Correlated Linearizations and Test Error of Random Features Regression

### v2

@@ 6 Proof of Asymptotic Equivalence for Random Features
$\mathcal{S}:M:\mathbb{C}^{\ell\times\ell}\mapsto\begin{bmatrix}\mathrm{tr}(M_{2,2})K_{AA^{T}}&0&0&\mathrm{tr}(M_{2,2})K_{A\hat{A}^{T}}\\ 0&\rho(M)I_{d}&0&0\\ 0&0&0&0\\ \mathrm{tr}(M_{2,2})K_{\hat{A}A^{T}}&0&0&\mathrm{tr}(M_{2,2})K_{\hat{A}\hat{A}^{T}}\end{bmatrix}\in\mathbb{C}^{\ell\times\ell}$
where $\rho(M):=\mathrm{tr}(K_{AA^{T}}M_{1,1}+K_{A\hat{A}^{T}}M_{4,1}+K_{\hat{A}A^{T}}M_{1,4}+K_{\hat{A}\hat{A}^{T}}M_{4,4})$. Then, $\mathcal{S}(M)=\mathbb{E}[(L-\mathbb{E}L)M(L-\mathbb{E}L)]-\tilde{\mathcal{S}}(M)$ holds with
$\tilde{\mathcal{S}}(M):=\mathbb{E}\begin{bmatrix}0&\begin{subarray}{c}K_{AA^{T}}M^{T}_{2,1}+K_{A\hat{A}^{T}}M^{T}_{2,4}\end{subarray}&0&0\\ \begin{subarray}{c}M^{T}_{1,2}K_{AA^{T}}+M^{T}_{4,2}K_{\hat{A}A^{T}}\end{subarray}&0&0&\begin{subarray}{c}M_{1,2}^{T}K_{A\hat{A}^{T}}+M_{4,2}^{T}K_{\hat{A}\hat{A}^{T}}\end{subarray}\\ 0&0&0&0\\ 0&\begin{subarray}{c}K_{\hat{A}A^{T}}M^{T}_{2,1}+K_{\hat{A}\hat{A}^{T}}M^{T}_{2,4}\end{subarray}&0&0\end{bmatrix}.$
By Theorem 2.1, there exists a unique solution $M\in\mathcal{M}$ such that $M(z)$ solves (2) for every $z\in\mathbb{H}$. Plugging-in the expression for the superoperator above and using block inversion lemma, we find that
$M(z)=\begin{bmatrix}\begin{subarray}{c}((\delta-z)I_{n_{\text{train}}}-\mathrm{tr}(M_{2,2})K_{AA^{T}})^{-1}\end{subarray}&0&-\begin{subarray}{c}\mathrm{tr}(M_{2,2})M_{1,1}K_{A\hat{A}^{T}}\end{subarray}&0\\ 0&\begin{subarray}{c}d^{-1}\mathrm{tr}(M_{2,2})I_{d}\end{subarray}&0&0\\ \begin{subarray}{c}-\mathrm{tr}(M_{2,2})K_{\hat{A}A^{T}}M_{1,1}\end{subarray}&0&\begin{subarray}{c}(\mathrm{tr}(M_{2,2}))^{2}K_{\hat{A}A^{T}}M_{1,1}K_{A\hat{A}^{T}}\\ +\mathrm{tr}(M_{2,2})K_{\hat{A}\hat{A}^{T}}\end{subarray}&\begin{subarray}{c}-I_{n_{\text{test}}}\end{subarray}\\ 0&0&\begin{subarray}{c}-I_{n_{\text{test}}}\end{subarray}&0\end{bmatrix}$ (23)
with $M_{2,2}=-(1+z+\mathrm{tr}(K_{AA^{T}}M_{1,1}))^{-1}I_{d}$.
@@
$\mathcal{S}^{(\mathrm{sub})}:X\mapsto\begin{bmatrix}\mathrm{tr}(X_{2,2})K_{AA^{T}}&0\\ 0&\mathrm{tr}(K_{AA^{T}}X_{1,1})\end{bmatrix}$
@@
Here, the set +$\mathcal{M}^{(\mathrm{sub})}_{+}=\mathrm{Hol}(\mathbb{H},\mathcal{A}^{(\mathrm{sub})})$ and $\mathcal{A}^{(\mathrm{sub})}=\{N\in\mathbb{C}^{(n_{\text{train}}+d)\times(n_{\text{train}}+d)}\,:\,\Im[N]\succ 0\}$. Given that the sub-DEL has a spectral parameter spanning its diagonal, the iteration scheme $N_{k+1}=\mathcal{F}^{(\mathrm{sub})}(N_{k})$ converges to the unique solution of the sub-DEL $M^{(\mathrm{sub})}=\mathcal{F}^{(\mathrm{sub})}(M^{(\mathrm{sub})})$ for any $N_{0}\in\mathcal{M}^{(\mathrm{sub})}_{+}$, as per Proposition 4.15.
@@
$|\|n^{-\frac{1}{2}}\sigma(XW_{1})\|-\|n^{-\frac{1}{2}}\sigma(XW_{2})\||$ $\leq|n^{-\frac{1}{2}}\sigma(XW_{1})-n^{-\frac{1}{2}}\sigma(XW_{2})\|$
@@
By Gaussian concentration inequality for Lipschitz functions [58, 44], there exists absolute constants $c_{1},c_{2}\in\mathbb{R}_{>0}$ such that $\mathbb{P}(|\|A\|-\mathbb{E}\|A\||\geq t)\leq c_{1}\exp(-c_{2}nt^{2}/(\lambda_{\sigma}^{2}\lambda_{\varphi}^{2}\|X\|^{2}))$. In other words, the random variable $\|A\|-\mathbb{E}\|A\|$ is $\lambda_{\sigma}\lambda_{\varphi}\|X\|/\sqrt{n}$-sub-Gaussian, and similarly for $\|\hat{A}\|-\mathbb{E}\|\hat{A}\|$. The result follows by applying the moment bound for sub-Gaussian random variables and using the fact that $\mathbb{E}\|A\|$ and $\mathbb{E}\|\hat{A}\|$ are bounded as $n\to\infty$. ∎
In the following sections, we use the Dyson equation for linearization framework to derive a deterministic equivalent for the matrix $\hat{A}A^{T}(AA^{T}+\delta I_{n_{\text{train}}})^{-1}$ in Section 6.1, and subsequently derive a deterministic equivalent for its square in Section 6.2.
@@ 6.1 First Deterministic Equivalent
$\Re[\mathcal{F}^{(\mathrm{sub})}(N)]=\mathcal{F}^{(\mathrm{sub})}(N)\begin{bmatrix}\begin{subarray}{c}(\delta-\Re[z])I_{n_{\text{train}}}\\ -\mathrm{tr}(\Re[N_{2,2}])K_{AA^{T}}\end{subarray}&0\\ 0&\begin{subarray}{c}-(1+\Re[z]+\mathrm{tr}(K_{AA^{T}}\Re[N_{1,1}]))\end{subarray}\end{bmatrix}(\mathcal{F}^{(\mathrm{sub})}(N))^{\ast}$ (24)
@@
We first derive some useful norm bounds. Recall that $R=((1+z)^{-1}AA^{T}+(\delta-z)I_{n_{\text{train}}})^{-1}$. For $|z|<1\wedge\delta$, $\Re[(1+z)^{-1}]\geq|1+z|^{-2}(1-|z|)\geq(1-|z|)/4>0$ and $\Re[\delta-z]\geq\delta-|z|$. Hence, $\Re[R]\geq(\delta-|z|)RR^{\ast}$ which implies that $\|R\|\leq(\delta-|z|)^{-1}$. A similar argument applied to $\bar{R}=-((1+z)I_{d}+(\delta-z)^{-1}A^{T}A)^{-1}$ gives $\|\bar{R}\|\leq(1-|z|)^{-1}$. Furthermore, we know that $\|RA\|^{2}=\|RAA^{T}R^{\ast}\|\leq\|RAA^{T}\|\|R\|$. By definition, $RAA^{T}=(1+z)I_{n_{\text{train}}}-(1+z)(\delta-z)R$. Thus, $\|RAA^{T}\|\leq 2(2+\delta)(\delta-|z|)^{-1}$ and $\|RA\|\leq\sqrt{2(2+\delta)}(\delta-|z|)^{-1}$.
We now check that $L^{(\mathrm{sub})}$ satisfies conditions (C3), (C1) and (C2). Since $R$ and $\bar{R}$ as well as the moments of $\|A\|,\|\hat{A}\|$ are bounded as $n$ increases, $\limsup_{\ell\to\infty}\mathbb{E}\|(L-z\Lambda)^{-1}\|^{2}<\infty$. In particular, condition (C1) is satisfied. Since there is a spectral parameter spanning the entire diagonal in the sub-DEL, it follows from the stability properties of the DEL or [10, Corollary 3.8] that condition (C2) is satisfied for the sub-DEL. It follows from Lemma 6.2 that $\limsup_{\ell\to\infty}(\|\mathcal{S}\|\vee\|\mathbb{E}L\|)<\infty$. Finally, condition (C3) is evidently satisfied. Hence, the conditions are satisfied.
@@
$\|\mathcal{S}^{(\mathrm{sub})}(N)\|\leq\|K_{AA^{T}}\||\mathrm{tr}(N_{2,2})|+|\mathrm{tr}(K_{AA^{T}}N_{1,1})|\leq(\sqrt{d}+\sqrt{n_{\text{train}}})\|K_{AA^{T}}\|\|N\|_{F}.$
By Jensen’s inequality, $\|K_{AA^{T}}\|=\|d^{-1}\mathbb{E}[AA^{T}]\|\leq d^{-1}\mathbb{E}\|A\|^{2}$. In fact, by a similar argument, $\|K_{AA^{T}}\|\vee\|K_{\hat{A}A^{T}}\|\vee\|K_{A\hat{A}^{T}}\|\vee\|K_{\hat{A}\hat{A}^{T}}\|\lesssim n^{-1}$. Since $\mathbb{E}\|A\|$ is bounded by Lemma 6.2, and we are working in the proportional limit, $\|\mathcal{S}^{(\mathrm{sub})}\|_{F\mapsto 2}\lesssim n^{-1/2}$. Next, let $Z_{1},Z_{2}\in\mathbb{R}^{n_{0}\times d}$ and notice that $\|\mathcal{C}(Z_{1})-\mathcal{C}(Z_{2})\|_{F}\leq n^{-1/2}\lambda_{\sigma}\lambda_{\varphi}\|X\|\|Z_{1}-Z_{2}\|_{F}$. Since $\lambda_{\sigma}$, $\lambda_{\varphi}$ and $\|X\|$ are all bounded by assumption, $\lambda\lesssim(\Im[z])^{-2}n^{-1}$. Finally, for every $N\in\mathbb{C}^{(n_{\text{train}}+d)\times(n_{\text{train}}+d)}$,
$\tilde{\mathcal{S}}^{(\mathrm{sub})}(N)=\mathbb{E}\begin{bmatrix}0&K_{AA^{T}}N^{T}_{2,1}+K_{A\hat{A}^{T}}N^{T}_{2,4}\\ N^{T}_{1,2}K_{AA^{T}}+N^{T}_{4,2}K_{\hat{A}A^{T}}&0\end{bmatrix}.$
Since $\|K_{AA^{T}}\|\vee\|K_{\hat{A}A^{T}}\|\vee\|K_{A\hat{A}^{T}}\|\vee\|K_{\hat{A}\hat{A}^{T}}\|\lesssim n^{-1}$, $\|\tilde{\mathcal{S}}^{(\mathrm{sub})}\|\lesssim n^{-1}$. Combining everything along with concentration of linear functional of the resolvent around their mean, the result follows from Theorem 2.3. ∎
@@
Fix $z\in\mathbb{H}$ with $|z|<1\wedge\delta$ and let $\tau\in\mathbb{R}_{>0}$. Expanding (5), we have $((\delta-z-i\tau)I_{n_{\text{train}}}-\mathrm{tr}(M_{2,2}^{(\tau)})K_{AA^{T}})M_{1,4}^{(\tau)}=\mathrm{tr}(M_{2,2}^{(\tau)})K_{A\hat{A}^{T}}M_{4,4}^{(\tau)}$, $-i\tau M_{3,4}^{(\tau)}-M_{4,4}^{(\tau)}=0$ and $-M_{3,4}^{(\tau)}-(\mathrm{tr}(M_{2,2}^{(\tau)})K_{\hat{A}\hat{A}^{T}}+i\tau I_{n_{\text{train}}})M_{4,4}^{(\tau)}-\mathrm{tr}(M_{2,2}^{(\tau)})K_{\hat{A}A^{T}}M_{1,4}^{(\tau)}=I_{n_{\text{test}}}$. Using those equations, we may solve for $M_{4,4}^{(\tau)}$ and $M_{1,4}^{(\tau)}$. In particular, we have
$M_{1,4}^{(\tau)}=\mathrm{tr}(M_{2,2}^{(\tau)})((\delta-z-i\tau)I_{n_{\text{train}}}-\mathrm{tr}(M_{2,2}^{(\tau)})K_{AA^{T}})^{-1}K_{A\hat{A}^{T}}M_{4,4}^{(\tau)}$
and $(I_{n_{\text{test}}}-i\tau\Xi)M_{4,4}^{(\tau)}=i\tau I_{n_{\text{test}}}$ with $\Xi=(\mathrm{tr}(M_{2,2}^{(\tau)})K_{\hat{A}\hat{A}^{T}}+i\tau I_{n_{\text{train}}})-(\mathrm{tr}(M_{2,2}^{(\tau)}))^{2}K_{\hat{A}A^{T}}((\delta-z-i\tau)I_{n_{\text{train}}}-\mathrm{tr}(M_{2,2}^{(\tau)})K_{AA^{T}})^{-1}K_{A\hat{A}^{T}}$. Since $\|M_{2,2}^{(\tau)}\|\leq(\Im[z])^{-1}$ and $\Im[M^{(\tau)}_{2,2}]\succeq 0$ for every $\tau\in\mathbb{R}_{>0}$, it follows that $\|\Xi\|$ is bounded as $\tau\to 0$. Consequently, for every $\tau$ small enough, $\|i\tau\Xi\|<1$ and $I_{n_{\text{test}}}-i\tau\Xi$ is invertible. In particular, using the fact that $\|K_{AA^{T}}\|\vee\|K_{\hat{A}A^{T}}\|\vee\|K_{A\hat{A}^{T}}\|\vee\|K_{\hat{A}\hat{A}^{T}}\|\lesssim n^{-1}$, $M_{4,4}^{(\tau)}$ approaches 00 as $\tau\to 0$ uniformly in $n$. The same argument can be applied to $M_{1,4}^{(\tau)}$.
@@
$D^{(\mathrm{sub})}=\mathrm{diag}\{\mathrm{tr}(M_{2,2}^{(\tau)})K_{A\hat{A}^{T}}M_{4,1}^{(\tau)},\mathrm{tr}(K_{A\hat{A}^{T}}M^{(\tau)}_{4,1}+K_{\hat{A}A^{T}}M^{(\tau)}_{1,4}+K_{\hat{A}\hat{A}^{T}}M^{(\tau)}_{4,4})M_{2,2}^{(\tau)}\}.$
@@
Having established that the solution to (23) acts as a deterministic equivalent for the pseudo-resolvent $(L-z\Lambda)^{-1}$ linked to (22), we aim to retrieve the expression in (11) by taking the spectral parameter to 00. To accomplish this, we need further control over the DEL near the origin.
@@
Let $z\in\mathbb{H}$ with $|z|<\delta\wedge 1$ and $M\in\mathcal{M}$ the unique solution to (23). Then, for every $\epsilon\in(0,2^{-1}]$ with $(2(\delta-\Re[z])^{-2}d^{2}\|K_{AA^{T}}\|^{2}-1-\Re[z])\epsilon\leq 2^{-1}(1+\Re[z])$,
$1-d\|K_{AA^{T}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{T}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}(z)\|^{2}\geq\epsilon.$
@@
$\limsup_{n\to\infty}d\|K_{AA^{T}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{T}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}(z)\|^{2}<1-\epsilon$

### v4

@@ 6 Proof of Asymptotic Equivalence for Random Features
$\mathcal{S}:M\in\mathbb{C}^{\ell\times\ell}\mapsto\begin{bmatrix}\mathrm{tr}(M_{2,2})K_{AA^{\top}}&0&0&\mathrm{tr}(M_{2,2})K_{A\hat{A}^{\top}}\\ 0&\rho(M)I_{d}&0&0\\ 0&0&0&0\\ \mathrm{tr}(M_{2,2})K_{\hat{A}A^{\top}}&0&0&\mathrm{tr}(M_{2,2})K_{\hat{A}\hat{A}^{\top}}\end{bmatrix}\in\mathbb{C}^{\ell\times\ell}$
where $\rho(M):=\mathrm{tr}(K_{AA^{\top}}M_{1,1}+K_{A\hat{A}^{\top}}M_{4,1}+K_{\hat{A}A^{\top}}M_{1,4}+K_{\hat{A}\hat{A}^{\top}}M_{4,4})$. Then, $\mathcal{S}(M)=\mathbb{E}[(L-\mathbb{E}L)M(L-\mathbb{E}L)]-\tilde{\mathcal{S}}(M)$ holds with
$\tilde{\mathcal{S}}(M):=\mathbb{E}\begin{bmatrix}0&\begin{subarray}{c}K_{AA^{\top}}M^{\top}_{2,1}+K_{A\hat{A}^{\top}}M^{\top}_{2,4}\end{subarray}&0&0\\ \begin{subarray}{c}M^{\top}_{1,2}K_{AA^{\top}}+M^{\top}_{4,2}K_{\hat{A}A^{\top}}\end{subarray}&0&0&\begin{subarray}{c}M_{1,2}^{\top}K_{A\hat{A}^{\top}}+M_{4,2}^{\top}K_{\hat{A}\hat{A}^{\top}}\end{subarray}\\ 0&0&0&0\\ 0&\begin{subarray}{c}K_{\hat{A}A^{\top}}M^{\top}_{2,1}+K_{\hat{A}\hat{A}^{\top}}M^{\top}_{2,4}\end{subarray}&0&0\end{bmatrix}$.
By Theorem 2.1, there exists a unique solution $M\in\mathcal{M}$ such that $M(z)$ solves (2) for every $z\in\mathbb{H}$. Plugging-in the expression for the superoperator above and using Lemma A.9, we find that
$M(z)=\begin{bmatrix}\begin{subarray}{c}((\delta-z)I_{n_{\text{train}}}-\mathrm{tr}(M_{2,2})K_{AA^{\top}})^{-1}\end{subarray}&0&-\begin{subarray}{c}\mathrm{tr}(M_{2,2})M_{1,1}K_{A\hat{A}^{\top}}\end{subarray}&0\\ 0&\begin{subarray}{c}d^{-1}\mathrm{tr}(M_{2,2})I_{d}\end{subarray}&0&0\\ \begin{subarray}{c}-\mathrm{tr}(M_{2,2})K_{\hat{A}A^{\top}}M_{1,1}\end{subarray}&0&\begin{subarray}{c}(\mathrm{tr}(M_{2,2}))^{2}K_{\hat{A}A^{\top}}M_{1,1}K_{A\hat{A}^{\top}}\\ +\mathrm{tr}(M_{2,2})K_{\hat{A}\hat{A}^{\top}}\end{subarray}&\begin{subarray}{c}-I_{n_{\text{test}}}\end{subarray}\\ 0&0&\begin{subarray}{c}-I_{n_{\text{test}}}\end{subarray}&0\end{bmatrix}$ (23)
with $M_{2,2}=-(1+z+\mathrm{tr}(K_{AA^{\top}}M_{1,1}))^{-1}I_{d}$.
@@
$\mathcal{S}^{(\mathrm{sub})}:X\mapsto\begin{bmatrix}\mathrm{tr}(X_{2,2})K_{AA^{\top}}&0\\ 0&\mathrm{tr}(K_{AA^{\top}}X_{1,1})\end{bmatrix}$
@@
Here, the set $\mathcal{M}^{(\mathrm{sub})}_{+}=\mathrm{Hol}(\mathbb{H},\mathcal{A}^{(\mathrm{sub})})$ and $\mathcal{A}^{(\mathrm{sub})}=\{N\in\mathbb{C}^{(n_{\text{train}}+d)\times(n_{\text{train}}+d)}\,:\,\Im[N]\succ 0\}$. Given that the sub-DEL has a spectral parameter spanning its diagonal, the iteration scheme $N_{k+1}=\mathcal{F}^{(\mathrm{sub})}(N_{k})$ converges to the unique solution of the sub-DEL $M^{(\mathrm{sub})}=\mathcal{F}^{(\mathrm{sub})}(M^{(\mathrm{sub})})$ for any $N_{0}\in\mathcal{M}^{(\mathrm{sub})}_{+}$, as per Proposition 4.1.
@@
[…]
@@
By Gaussian concentration inequality for Lipschitz functions [tao_rmt, ledoux_concentration_2001], there exists absolute constants $c_{1},c_{2}\in\mathbb{R}_{>0}$ such that $\mathbb{P}(|\|A\|-\mathbb{E}\|A\||\geq t)\leq c_{1}\exp(-c_{2}nt^{2}/(\lambda_{\sigma}^{2}\lambda_{\varphi}^{2}\|X\|^{2}))$. In other words, the random variable $\|A\|-\mathbb{E}\|A\|$ is $\lambda_{\sigma}\lambda_{\varphi}\|X\|/\sqrt{n}$-sub-Gaussian, and similarly for $\|\hat{A}\|-\mathbb{E}\|\hat{A}\|$. The result follows by applying the moment bound for sub-Gaussian random variables and using the fact that $\mathbb{E}\|A\|$ and $\mathbb{E}\|\hat{A}\|$ are bounded as $n\to\infty$. ∎
In the following sections, we use the Dyson equation for linearization framework to derive a deterministic equivalent for the matrix $\hat{A}A^{\top}(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}$ in Section 6.1, and subsequently derive a deterministic equivalent for its square in Section 6.2.
@@ 6.1 First Deterministic Equivalent
$\Re[\mathcal{F}^{(\mathrm{sub})}(N)]=\mathcal{F}^{(\mathrm{sub})}(N)\begin{bmatrix}\begin{subarray}{c}(\delta-\Re[z])I_{n_{\text{train}}}\\ -\mathrm{tr}(\Re[N_{2,2}])K_{AA^{\top}}\end{subarray}&0\\ 0&\begin{subarray}{c}-(1+\Re[z]+\mathrm{tr}(K_{AA^{\top}}\Re[N_{1,1}]))\end{subarray}\end{bmatrix}(\mathcal{F}^{(\mathrm{sub})}(N))^{\ast}$ (24)
@@
We first derive some useful norm bounds. Recall that $R=((1+z)^{-1}AA^{\top}+(\delta-z)I_{n_{\text{train}}})^{-1}$. For $|z|<1\wedge\delta$, $\Re[(1+z)^{-1}]\geq|1+z|^{-2}(1-|z|)\geq(1-|z|)/4>0$ and $\Re[\delta-z]\geq\delta-|z|$. Hence, $\Re[R]\geq(\delta-|z|)RR^{\ast}$ which implies that $\|R\|\leq(\delta-|z|)^{-1}$. A similar argument applied to $\bar{R}=-((1+z)I_{d}+(\delta-z)^{-1}A^{\top}A)^{-1}$ gives $\|\bar{R}\|\leq(1-|z|)^{-1}$. Furthermore, we know that $\|RA\|^{2}=\|RAA^{\top}R^{\ast}\|\leq\|RAA^{\top}\|\|R\|$. By definition, $RAA^{\top}=(1+z)I_{n_{\text{train}}}-(1+z)(\delta-z)R$. Thus, $\|RAA^{\top}\|\leq 2(2+\delta)(\delta-|z|)^{-1}$ and $\|RA\|\leq\sqrt{2(2+\delta)}(\delta-|z|)^{-1}$.
We now check that $L^{(\mathrm{sub})}$ satisfies conditions (C3), (C1) and (C2). Since $R$ and $\bar{R}$ as well as the moments of $\|A\|,\|\hat{A}\|$ are bounded as $n$ increases, $\limsup_{\ell\to\infty}\mathbb{E}\|(L-z\Lambda)^{-1}\|^{2}<\infty$. In particular, condition (C1) is satisfied. Since there is a spectral parameter spanning the entire diagonal in the sub-DEL, it follows from the stability properties of the DEL or [alt_Kronecker, Corollary 3.8] that condition (C2) is satisfied for the sub-DEL. It follows from Lemma 6.2 that $\limsup_{\ell\to\infty}(\|\mathcal{S}\|\vee\|\mathbb{E}L\|)<\infty$. Finally, condition (C3) is evidently satisfied. Hence, the conditions are satisfied.
@@
$\|\mathcal{S}^{(\mathrm{sub})}(N)\|\leq\|K_{AA^{\top}}\||\mathrm{tr}(N_{2,2})|+|\mathrm{tr}(K_{AA^{\top}}N_{1,1})|\leq(\sqrt{d}+\sqrt{n_{\text{train}}})\|K_{AA^{\top}}\|\|N\|_{F}.$
By Jensen’s inequality, $\|K_{AA^{\top}}\|=\|d^{-1}\mathbb{E}[AA^{\top}]\|\leq d^{-1}\mathbb{E}\|A\|^{2}$. In fact, by a similar argument, $\|K_{AA^{\top}}\|\vee\|K_{\hat{A}A^{\top}}\|\vee\|K_{A\hat{A}^{\top}}\|\vee\|K_{\hat{A}\hat{A}^{\top}}\|\lesssim n^{-1}$. Since $\mathbb{E}\|A\|$ is bounded by Lemma 6.2, and we are working in the proportional limit, $\|\mathcal{S}^{(\mathrm{sub})}\|_{F\mapsto 2}\lesssim n^{-1/2}$. Next, let $Z_{1},Z_{2}\in\mathbb{R}^{n_{0}\times d}$ and notice that $\|\mathcal{C}(Z_{1})-\mathcal{C}(Z_{2})\|_{F}\leq n^{-1/2}\lambda_{\sigma}\lambda_{\varphi}\|X\|\|Z_{1}-Z_{2}\|_{F}$. Since $\lambda_{\sigma}$, $\lambda_{\varphi}$ and $\|X\|$ are all bounded by assumption, $\lambda\lesssim(\Im[z])^{-2}n^{-1}$. Finally, for every $N\in\mathbb{C}^{(n_{\text{train}}+d)\times(n_{\text{train}}+d)}$,
$\tilde{\mathcal{S}}^{(\mathrm{sub})}(N)=\mathbb{E}\begin{bmatrix}0&K_{AA^{\top}}N^{\top}_{2,1}+K_{A\hat{A}^{\top}}N^{\top}_{2,4}\\ N^{\top}_{1,2}K_{AA^{\top}}+N^{\top}_{4,2}K_{\hat{A}A^{\top}}&0\end{bmatrix}$.
Since $\|K_{AA^{\top}}\|\vee\|K_{\hat{A}A^{\top}}\|\vee\|K_{A\hat{A}^{\top}}\|\vee\|K_{\hat{A}\hat{A}^{\top}}\|\lesssim n^{-1}$, $\|\tilde{\mathcal{S}}^{(\mathrm{sub})}\|\lesssim n^{-1}$. Combining everything along with concentration of linear functional of the resolvent around their mean, the result follows from Theorem 2.2. ∎
@@
Fix $z\in\mathbb{H}$ with $|z|<1\wedge\delta$ and let $\tau\in\mathbb{R}_{>0}$. Expanding (5), we have $((\delta-z-i\tau)I_{n_{\text{train}}}-\mathrm{tr}(M_{2,2}^{(\tau)})K_{AA^{\top}})M_{1,4}^{(\tau)}=\mathrm{tr}(M_{2,2}^{(\tau)})K_{A\hat{A}^{\top}}M_{4,4}^{(\tau)}$, $i\tau M_{3,4}^{(\tau)}-M_{4,4}^{(\tau)}=0$ and $M_{3,4}^{(\tau)}-(\mathrm{tr}(M_{2,2}^{(\tau)})K_{\hat{A}\hat{A}^{\top}}+i\tau I_{n_{\text{train}}})M_{4,4}^{(\tau)}-\mathrm{tr}(M_{2,2}^{(\tau)})K_{\hat{A}A^{\top}}M_{1,4}^{(\tau)}=I_{n_{\text{test}}}$. Using those equations, we may solve for $M_{4,4}^{(\tau)}$ and $M_{1,4}^{(\tau)}$. In particular, we have
$M_{1,4}^{(\tau)}=\mathrm{tr}(M_{2,2}^{(\tau)})((\delta-z-i\tau)I_{n_{\text{train}}}-\mathrm{tr}(M_{2,2}^{(\tau)})K_{AA^{\top}})^{-1}K_{A\hat{A}^{\top}}M_{4,4}^{(\tau)}$
and $(I_{n_{\text{test}}}-i\tau\Xi)M_{4,4}^{(\tau)}=i\tau I_{n_{\text{test}}}$ with $\Xi=(\mathrm{tr}(M_{2,2}^{(\tau)})K_{\hat{A}\hat{A}^{\top}}+i\tau I_{n_{\text{train}}})-(\mathrm{tr}(M_{2,2}^{(\tau)}))^{2}K_{\hat{A}A^{\top}}((\delta-z-i\tau)I_{n_{\text{train}}}-\mathrm{tr}(M_{2,2}^{(\tau)})K_{AA^{\top}})^{-1}K_{A\hat{A}^{\top}}$. Since $\|M_{2,2}^{(\tau)}\|\leq(\Im[z])^{-1}$ and $\Im[M^{(\tau)}_{2,2}]\succeq 0$ for every $\tau\in\mathbb{R}_{>0}$, it follows that $\|\Xi\|$ is bounded as $\tau\to 0$. Consequently, for every $\tau$ small enough, $\|i\tau\Xi\|<1$ and $I_{n_{\text{test}}}-i\tau\Xi$ is invertible. In particular, using the fact that $\|K_{AA^{\top}}\|\vee\|K_{\hat{A}A^{\top}}\|\vee\|K_{A\hat{A}^{\top}}\|\vee\|K_{\hat{A}\hat{A}^{\top}}\|\lesssim n^{-1}$, $M_{4,4}^{(\tau)}$ approaches 0 as $\tau\to 0$ uniformly in $n$. The same argument can be applied to $M_{1,4}^{(\tau)}$.
@@
$D^{(\mathrm{sub})}=\mathrm{diag}\{\mathrm{tr}(M_{2,2}^{(\tau)})K_{A\hat{A}^{\top}}M_{4,1}^{(\tau)},\mathrm{tr}(K_{A\hat{A}^{\top}}M^{(\tau)}_{4,1}+K_{\hat{A}A^{\top}}M^{(\tau)}_{1,4}+K_{\hat{A}\hat{A}^{\top}}M^{(\tau)}_{4,4})M_{2,2}^{(\tau)}\}.$
@@
Having established that the solution to (23) acts as a deterministic equivalent for the pseudo-resolvent $(L-z\Lambda)^{-1}$ linked to (22), we aim to retrieve the expression in (11) by taking the spectral parameter to 0. To accomplish this, we need further control over the DEL near the origin.
@@
Let $z\in\mathbb{H}$ with $|z|<\delta\wedge 1$ and $M\in\mathcal{M}$ the unique solution to (23). Then, for every $\epsilon\in(0,2^{-1}]$ with $(2(\delta-\Re[z])^{-2}d^{2}\|K_{AA^{\top}}\|^{2}-1-\Re[z])\epsilon\leq 2^{-1}(1+\Re[z])$,
$1-d\|K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{\top}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}(z)\|^{2}\geq\epsilon.$
@@
$\limsup_{n\to\infty}d\|K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{\top}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}(z)\|^{2}<1-\epsilon$